OpenStudy (anonymous):

How do I solve for x? 9|9 – 8x| = 2x + 3

OpenStudy (anonymous):

Simplifying x + 9 = 18 + -2x Reorder the terms: 9 + x = 18 + -2x Solving 9 + x = 18 + -2x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '2x' to each side of the equation. 9 + x + 2x = 18 + -2x + 2x Combine like terms: x + 2x = 3x 9 + 3x = 18 + -2x + 2x Combine like terms: -2x + 2x = 0 9 + 3x = 18 + 0 9 + 3x = 18 Add '-9' to each side of the equation. 9 + -9 + 3x = 18 + -9 Combine like terms: 9 + -9 = 0 0 + 3x = 18 + -9 3x = 18 + -9 Combine like terms: 18 + -9 = 9 3x = 9 Divide each side by '3'. x = 3 Simplifying x = 3